An advertisement claims that a centrifuge takes up only 0.127 m of bench space but can 

produce a radial acceleration of 3000g (it is simply 3000 times the value of 𝑔 = 9.8𝑚/𝑠

) at 

5000 rev/min. Calculate the required radius of the centrifuge. Is the claim realistic?


1
Expert's answer
2021-06-14T15:03:33-0400

Gives


"a_c=3000\times g=3000\times9.8=29400m/sec^2"

"w=2\pi\ \times 5000\times \frac{1}{60}=523.33 rad/sec"

"a_c=rw^2"

Put values

"r=\frac {a_c}{w^2}"

"r=\frac{29400}{523.33^2}=0.107meter"

it the claim is not realistic